OpenStudy (anonymous):

x^2-3x+5=x(x-8)+20 help!!!

OpenStudy (anonymous):

x^2-3x+5=x^2-8x +20 -3x+5= -8x+20 5x=15 x=3

OpenStudy (anonymous):

thank you!!!

OpenStudy (ajprincess):

x^2-3x+5=x^2-8x+20 8x-3x=20-5 5x=15 x=3

OpenStudy (anonymous):

thank you!!

OpenStudy (ajprincess):

yw